OpenStudy (anonymous):

Solve: -5 + x/3 = 20 can you lease do it and show all of the steps because I barly have enough time to turn it in thanks

OpenStudy (anonymous):

-5 + x/3 = 20 first step is to isolate the x term by moving the 5 over to the 20 side x/3 = 20+5 x/3 = 25 next step is to multiply both sides by 3 to get rid of the fraction that the x is in 3(x/3) = (25)3 x = 75 make sense?